a man is ten times as old as his daughter, in 5 years time he will be 5 times as old as his daughter. how old are each of them now ?
Please send breakdown of solution
a father is 48 years old and his son 14 years old ,in how many years will the father be three times older than the son?
Please send break down and solution

Answer by mananth(16946)

daughter 's age =x
father's age = 10x
5 years later
daughter's age = x+5
father's age = 10x+5
10x+5 = 5(x+5)
10x+5=5x+25
10x-5x=25-5
5x=20
x=4
daughter 4 years
Father 40 years